Show that the square of an even number is an even number using a direct proof.


1
Expert's answer
2022-06-15T05:52:19-0400

Let "n\in\mathbb{Z}" - even number, so "n=2z", where "z\in\mathbb{Z}".


"n^2=(2k)^2=2^2k^2=2\cdot(2k^2)" - is an even number because it is divisible by 2.